Assume a project has a sales quantity of 9,000 units, plus or minus 5 percent and a sales price of $80 a unit, plus or minus 5 percent. The expected variable cost per unit is $20, ±5 percent and the expected fixed costs are $310,000plus or minus 2 percent. The depreciation expense is $78,000. The tax rate is 34 percent. What is the operating cash flow under the best-case scenario?
Answer #1

Sales 793,800 =9000*(1+5%)*80*(1+5%) =9000*(1+5%)*20*(1-5%)+310000*(1-2%) Less: Costs Depreciation EBIT Less: Tax payable @ 3

*Please rate thumbs up